(6 points) The continuous random variable X has cumulative distribution function given by 0 for x <0 for x > 2 Part(a) Find the value of c correct to one decimal place given that E(X c) 4E(X - c). 0.4 Part(b) Two independent observations of X are taken. Find the probability correct to 2 decimal places that one is less than 1 and the other is greater than 1. The order in which we take the observations matters. Part(c) Find E(VX), correct to 2 decimal places. Part(d) Find E(X), correct to 2 decimal places. Part(e) Find Frx > 하 Give your answer as a decimal, correct to 2 decimal places. Part(f) Find the median of X, correct to 2 decimal places.
Answer #1

Solutian: Let X be 7v. with pdf 2 E (x+C)= 4Cr-c) E (x)+C=Ex -4C 3 Ex)= SC pdf of x fa) Fr fen)= 1-3 1-32x ECx) Saufin d =D-